OpenStudy (anonymous):

I need some things graphed. Can anyone help ? *note* looks long, but I'm jus including all info.

OpenStudy (anonymous):

Part 1: motor A: y= -16x^2 +190x+s a= -16 b= 190 c= 0 motor B y= -16x^2 +210x +s a= -16 b= 210 c= 0 motor c y= -16x^2 +230x+s a= -16 b= 230 c=0 motor D y= -16x^2 +250x+s a= -16 b=250 c= 0 Part 3: motor A (5.9, 564.04) motor B (6.5, 689) motor C ( 7.2, 826.56) motor D (7.8, 976.56) Part 4 13.1 seconds to return to the ground

OpenStudy (anonymous):

@amistre64 @AravindG @nubeer @satellite73 @Yahoo! can anyone help ????????

OpenStudy (nubeer):

umm is that 'S' in the equation..?

OpenStudy (anonymous):

I think so. This is what my partner gave me.

OpenStudy (nubeer):

hmm i think it should be some number not an alphabet.. and do u have to plot these on graph?

OpenStudy (anonymous):

yes I need them graphed. and I'm not too sure what it would be replaced with, it confused me.

OpenStudy (nubeer):

try this link to graph... http://my.hrw.com/math06_07/nsmedia/tools/Graph_Calculator/graphCalc.html normally just put in equations it shows u how ur graph will be.. and i am not still sure what that s in the end of every question is for.. maybe ask ur friend again if that is or just a typo

OpenStudy (anonymous):

ok thanks :)

OpenStudy (nubeer):

you're welcome.. :)